5) Box 1 contains wi white balls and bi black balls. Box 2 contains w2 white balls and b2 black balls. We take one ball from Box 1 and place it into Box 2. Then we take a ball from Box 2 and place it into Box 1. Finally we take a ball from Box 1. Compute the probability that this ball is black.
Answer #1

This can happen in 4 ways.
(a) placing white ball from Box 1 to Box 2 and then placing a white ball from Box 2 to Box 1.
Prob.(black ball from Box 1) = b1/(b1 + w1).

(b) placing white ball from Box 1 to Box 2 and then placing a black ball from Box 2 to Box 1.
Prob.(black ball from Box 1) = (b1 + 1)/(b1 + w1).

(c) placing black ball from Box 1 to Box 2 and then placing a white ball from Box 2 to Box 1.
Prob.(black ball from Box 1) = (b1 - 1)/(b1 + w1).

(d) placing black ball from Box 1 to Box 2 and then placing a black ball from Box 2 to Box 1.
Prob.(black ball from Box 1) = b1/(b1 + w1).

Total probability = adding the above 4 probabilities = (4 * b1)/(b1 + w1).
.